# Retrieval-Augmented Generation for Knowledge-Intensive NLP Tasks

### 22 May 2020 v1

@@ Abstract
Large pre-trained language models have been shown to store factual knowledge in their parameters, and achieve state-of-the-art results when fine-tuned on downstream NLP tasks. However, their ability to access and precisely manipulate knowledge is still limited, and hence on knowledge-intensive tasks, their performance lags behind task-specific architectures. Additionally, providing provenance for their decisions and updating their world knowledge remain open research problems. Pre-trained models with a differentiable access mechanism to explicit non-parametric memory can overcome this issue, but have so far been only investigated for extractive downstream tasks. We explore a general-purpose fine-tuning recipe for retrieval-augmented generation (RAG) — models which combine pre-trained parametric and non-parametric memory for language generation. We introduce RAG models where the parametric memory is a pre-trained seq2seq model and the non-parametric memory is a dense vector index of Wikipedia, accessed with a pre-trained neural retriever. We compare two RAG formulations, one which conditions on the same retrieved passages across the whole generated sequence, the other can use different passages per token. We fine-tune and evaluate our models on a wide range of knowledge-intensive NLP tasks and set the state-of-the-art on three open domain QA tasks, outperforming parametric seq2seq models and task-specific retrieve-and-extract architectures. For language generation tasks, we find that RAG models generate more specific, diverse and factual language than a state-of-the-art parametric-only seq2seq baseline.
@@ 1 Introduction
Pre-trained neural language models have been shown to learn a substantial amount of in-depth knowledge from data [41]. They can do so without any access to an external memory, as a parameterized implicit knowledge base [45, 46]. While this development is exciting, such models do have downsides: They cannot easily expand or revise their memory, can’t straightforwardly provide insight into their predictions, and may produce “hallucinations” [34]. Hybrid models that combine parametric memory with non-parametric (i.e., retrieval-based) memories [18, 22, 42] can address some of these issues because knowledge can be directly revised and expanded, and its access can be inspected and interpreted. REALM [18] and ORQA [27], two recently introduced models that combine masked language models [8] with a differentiable retriever, have shown promising results, but have only explored open-domain extractive question answering. Here, we bring hybrid parametric and non-parametric memory to the “workhorse of NLP,” i.e. sequence-to-sequence (seq2seq) models.
@@
We endow pre-trained, parametric-memory generation models with a non-parametric memory through a general-purpose fine-tuning approach which we refer to as retrieval-augmented generation (RAG). We build RAG models where the parametric memory is a pre-trained generative seq2seq transformer, and the non-parametric memory is a dense vector index of Wikipedia, accessed using a pre-trained neural retriever. We combine these components in an end-to-end probabilistic model; the document retriever (Dense Passage Retriever [22], henceforth DPR) provides latent documents conditioned on the input, and the seq2seq model (BART [28]) then conditions on both these latent documents and the input to generate the output. We marginalize the latent variables through a top-K approximation, either on a per answer basis (assuming the same document is responsible for all tokens) or a per answer token basis (assuming different documents can be responsible for different tokens). Just like T5 [45] or BART, RAG can be fine-tuned on any seq2seq task, whereby both the sequence generator and retriever are jointly learned.
@@
There has been extensive previous work proposing architectures to enrich systems with non-parametric memory which are trained from scratch for specific tasks—e.g. in memory networks [58, 49], stack-augmented networks [21] and memory layers for transformers [26]. In contrast, we explore a setting where both parametric and non-parametric memory components are pre-trained and pre-loaded with extensive knowledge. Crucially, by using pre-trained knowledge-access mechanisms, the ability to access knowledge is present without additional training.
@@
Our results highlight the benefits of combining parametric and non-parametric memory with generation for knowledge-intensive tasks. Our RAG models achieve state-of-the-art results on open Natural Questions [25], WebQuestions [3] and CuratedTrec [2] and strongly outperform recent approaches that use specialised pre-training objectives on TriviaQA [20]. Despite these being extractive tasks, we find that unconstrained generation outperforms previous extractive approaches. For knowledge-intensive generation, we experiment with MS-MARCO [1] and Jeopardy question generation, and we find that our models generate responses that are more factual, specific, and diverse than a BART baseline. For the FEVER [50] fact verification task, we achieve results within 4% of sophisticated, state-of-the-art pipeline models which use strong supervision. Finally, we show that the non-parametric memory can be replaced in order to control generation, demonstrating a simple mechanism to update the knowledge that the model uses as facts about the world change.
@@
Figure 1: An overview of retrieval-augmented generation (RAG). We combine a pre-trained retriever (Query Encoder + Document Index) with a pre-trained encoder-decoder (Generator) and fine-tune end-to-end. For some query $x$, we use Maximum Inner Product Search (MIPS) to find the top-K most relevant documents of all documents $z_{i}$. To make the final prediction $y$, we treat $z$ as a latent variable and marginalize over the encoder-decoder predictions given different documents.
@@ 2 Methods
We explore RAG models which use the input sequence $x$ to retrieve text passages $z$ and use these passages as additional context when generating the target sequence $y$. As shown in Figure 1, our models leverage two components: (i) a retriever $p_{\eta}(z|x)$ with parameters $\eta$ that returns (top-K truncated) distributions over text passages given a query $x$ and (ii) a generator $p_{\theta}(y_{i}|x,z,y_{1:i-1})$ parametrized by $\theta$ that generates a current token based on a context of the previous i1i-1 tokens $y_{1:i-1}$, the original input $x$ and a retrieved passage $z$.
To train the retriever and generator end-to-end, we treat the retrieved document as a latent variable. We propose two models that marginalize over the latent documents in different ways to produce a distribution over generated text. In one approach, RAG-Sequence, the model uses the same document to predict each target token. In the other approach, RAG-Token, the model can predict each target token based on a different document. In what follows, we formally introduce both models and then describe the $p_{\eta}$ and $p_{\theta}$ components, as well as the training and decoding procedure in more detail.
@@ 2.1 Models
The RAG-Sequence model uses the same retrieved document to generate the complete sequence. Technically, it treats the retrieved passage as a single latent variable that is marginalized to get the seq2seq probability $p(y|x)$ via a top-K approximation,
$p_{\text{\tiny{RAG-Sequence}}}(y|x)=\sum_{z\in\text{top-}k(p(\cdot|x))}p_{\eta}(z|x)\prod_{i}^{N}p_{\theta}(y_{i}|x,z,y_{1:i-1}).$
@@
In the RAG-Token model we can draw a different latent passage for each target token and marginalize accordingly. This allows the generator to choose content from several documents when producing an answer. Formally, we define:
$p_{\text{\tiny{RAG-Token}}}(y|x)=\prod_{i}^{N}\sum_{z\in\text{top-}k(p(\cdot|x))}p_{\eta}(z_{i}|x)p_{\theta}(y_{i}|x,z_{i},y_{1:i-1}).$
@@ 2.2 Retriever: DPR
$p_{\eta}(z|x)\propto\exp\left<\mathbf{d}(z),\mathbf{q}(x)\right$>
where $\mathbf{d}(z)$ is a dense representation of the document produced by a BERTBASE transformer [8], and $\mathbf{q}(x)$ a representation of the query by another BERTBASE transformer with a different set of parameters.
To efficiently calculate $\text{top-k}(p_{\eta}(\cdot|x))$, the list of $k$ elements $z$ with highest prior probability $p_{\eta}(z|x)$, DPR employs a Maximum Inner Product Search (MIPS) index provided by the FAISS library [19].
For non-parametric pre-trained memory, we use a pre-trained bi-encoder from [22] to both initialize our retriever and to build the document index. This retriever was trained to retrieve documents which contain answers to TriviaQA [20] questions and Natural Questions [25].
@@ 2.3 Generator: BART
The generator component $p_{\theta}(y_{i}|x,z,y_{1:i-1})$ could be modelled using any encoder-decoder. We use BART-large [28], a pre-trained seq2seq transformer [52] with 400M parameters. To combine the input $x$ with the retrieved content $z$ when generating from BART, we simply concatenate them.
BART was pre-trained using a denoising objective and a variety of different noising functions. It has obtained state-of-the-art results on a diverse set of generation tasks and outperforms comparably-sized T5 models [28]. We refer to the BART generator parameters $\theta$ as the parametric memory henceforth.
@@ 2.4 Training
We jointly train the retriever and generator components without any direct supervision on what document should be retrieved. Given a fine-tuning training corpus of input/output pairs $(x_{j},y_{j})$, we minimize the negative marginal log-likelihood of each target, $\sum_{j}-\log p(y_{j}|x_{j})$ using stochastic gradient descent with Adam [24]. Updating the document encoder during training is costly as it requires the document index to be periodically updated as REALM does during pre-training [18]. We do not find this step necessary for strong performance, and we keep the document encoder (and index) fixed, only fine-tuning the query encoder and the generator.
@@ 2.5 Decoding
At test/decoding time, RAG-Sequence and RAG-Token require different ways to approximate $\operatorname*{arg\,max}_{y}p(y|x)$.
@@
The RAG-Token model can be seen as a standard, autoregressive, seq2seq generator with transition probability:
$p^{\prime}_{\theta}(y_{i}|x,y_{1:i-1})=\sum_{z\in\text{top-}k(p(\cdot|x))}p_{\eta}(z_{i}|x)p_{\theta}(y_{i}|x,z_{i},y_{1:i-1})$
To decode, we can plug $p^{\prime}_{\theta}(y_{i}|x,y_{1:i-1})$ into a standard beam decoder.
@@
The likelihood $p(y|x)$ does not break into a conventional per-token likelihood for the RAG-Sequence, and hence we cannot solve it with a single beam search pass. Instead, we run beam search for each candidate document $z$, scoring each hypothesis using $p_{\theta}(y_{i}|x,z,y_{1:i-1})$. This yields a set of hypotheses $Y$ of which some might not have appeared in the beams of all documents. To estimate the probability of an hypothesis $y$ across all beams, we run an additional forward pass for each document $z$ for which $y$ does not appear in the beam, multiply the generator score with $p_{\eta}(z|x)$ and then sum up the probabilities across beams for the marginals. We refer to this decoding procedure as “Thorough Decoding.”
For longer output sequences, $|Y|$ can become large, requiring many forward passes. For more efficient decoding, we can make a further approximation that $p_{\theta}(y|x,z_{i})\approx 0$ where $y$ was not generated during beam search from $x,z_{i}$. This avoids the need to run additional forward passes once the candidate set $Y$ has been generated. We refer to this decoding procedure as “Fast Decoding”.
@@ 3 Experiments
We experiment with RAG in a wide range of knowledge-intensive tasks. For all experiments, we use a single Wikipedia dump for our non-parametric knowledge source. Following Lee et al. [27] and Karpukhin et al. [22], we use the December 2018 dump. Each Wikipedia article is split into disjoint 100-word chunks, to make a total of 21,015,324 documents. [^1] The reader is referred to Karpukhin et al. [22] for further details on how Wikipedia is pre-processed. We use the DPR document encoder to compute document embeddings for each document, and we build a single MIPS index using FAISS [19] using Hierarchical Navigable Small World approximation for efficient retrieval [33], which is then used for all experiments. During training, we retrieve the top $k$ documents for each query, where we consider $k\in\{5,10\}$. We determine $k$ for test time using validation data. In the remainder of this section, we will discuss the experimental details for each of these task settings.
@@ 3.1 Open-domain Question Answering
Open-domain QA is an important real-world NLP application and is often used as test-bed for knowledge-intensive tasks [18]. We tackle open-domain QA by treating questions and answers as simple input-output text pairs $(x,y)$, and we train RAG by directly minimizing the negative log-likelihood of answers. We compare our results to the popular extractive QA paradigm [5, 7, 27, 22], where answers are extracted as spans from retrieved documents, relying primarily on non-parametric knowledge. In addition, we also compare to "Closed-Book QA" approaches [46], which, like RAG, generate answers, but do not exploit latent retrieval, instead relying purely on parametric knowledge.
We consider four popular open-domain QA datasets: Natural Questions (NQ) [25], TriviaQA (TQA) [20]. WebQuestions (WQ) [3] and CuratedTrec (CT) [2]. The answers for CuratedTrec are given in the form of regular expressions, which has been cited as a reason why it is unsuitable for answer-generation models [18]. To overcome this, we use a pre-processing step where we first retrieve the top 1000 documents for each query, and use the answer that most frequently matches the regex pattern as the supervision target. If no matches are found, we resort to a simple heuristic: generate all possible permutations for each regex, replacing non-deterministic symbols in the regex nested tree structure with a whitespace. As CuratedTrec and WebQuestions are small datasets, we follow DPR [22] by initializing CuratedTrec and WebQuestions models with our Natural Questions RAG model.
We use the same training/dev/testing splitting method as in previous work [27, 22] and report the standard Exact Match (EM) metric. For TriviaQA, in order to compare to T5 [46], we do an additional test evaluation on the TriviaQA Wiki test set.

### 07 Dec 2020 v2

@@ Abstract
Large pre-trained language models have been shown to store factual knowledge in their parameters, and achieve state-of-the-art results when fine-tuned on downstream NLP tasks. However, their ability to access and precisely manipulate knowledge is still limited, and hence on knowledge-intensive tasks, their performance lags behind task-specific architectures. Additionally, providing provenance for their decisions and updating their world knowledge remain open research problems. Pre-trained models with a differentiable access mechanism to explicit non-parametric memory have so far been only investigated for extractive downstream tasks. We explore a general-purpose fine-tuning recipe for retrieval-augmented generation (RAG) — models which combine pre-trained parametric and non-parametric memory for language generation. We introduce RAG models where the parametric memory is a pre-trained seq2seq model and the non-parametric memory is a dense vector index of Wikipedia, accessed with a pre-trained neural retriever. We compare two RAG formulations, one which conditions on the same retrieved passages across the whole generated sequence, and another which can use different passages per token. We fine-tune and evaluate our models on a wide range of knowledge-intensive NLP tasks and set the state of the art on three open domain QA tasks, outperforming parametric seq2seq models and task-specific retrieve-and-extract architectures. For language generation tasks, we find that RAG models generate more specific, diverse and factual language than a state-of-the-art parametric-only seq2seq baseline.
@@ 1 Introduction
Pre-trained neural language models have been shown to learn a substantial amount of in-depth knowledge from data [47]. They can do so without any access to an external memory, as a parameterized implicit knowledge base [51, 52]. While this development is exciting, such models do have downsides: They cannot easily expand or revise their memory, can’t straightforwardly provide insight into their predictions, and may produce “hallucinations” [38]. Hybrid models that combine parametric memory with non-parametric (i.e., retrieval-based) memories [20, 26, 48] can address some of these issues because knowledge can be directly revised and expanded, and accessed knowledge can be inspected and interpreted. REALM [20] and ORQA [31], two recently introduced models that combine masked language models [8] with a differentiable retriever, have shown promising results, but have only explored open-domain extractive question answering. Here, we bring hybrid parametric and non-parametric memory to the “workhorse of NLP,” i.e. sequence-to-sequence (seq2seq) models.
@@
We endow pre-trained, parametric-memory generation models with a non-parametric memory through a general-purpose fine-tuning approach which we refer to as retrieval-augmented generation (RAG). We build RAG models where the parametric memory is a pre-trained seq2seq transformer, and the non-parametric memory is a dense vector index of Wikipedia, accessed with a pre-trained neural retriever. We combine these components in a probabilistic model trained end-to-end (Fig. 1). The retriever (Dense Passage Retriever [26], henceforth DPR) provides latent documents conditioned on the input, and the seq2seq model (BART [32]) then conditions on these latent documents together with the input to generate the output. We marginalize the latent documents with a top-K approximation, either on a per-output basis (assuming the same document is responsible for all tokens) or a per-token basis (where different documents are responsible for different tokens). Like T5 [51] or BART, RAG can be fine-tuned on any seq2seq task, whereby both the generator and retriever are jointly learned.
@@
There has been extensive previous work proposing architectures to enrich systems with non-parametric memory which are trained from scratch for specific tasks, e.g. memory networks [64, 55], stack-augmented networks [25] and memory layers [30]. In contrast, we explore a setting where both parametric and non-parametric memory components are pre-trained and pre-loaded with extensive knowledge. Crucially, by using pre-trained access mechanisms, the ability to access knowledge is present without additional training.
@@
Our results highlight the benefits of combining parametric and non-parametric memory with generation for knowledge-intensive tasks—tasks that humans could not reasonably be expected to perform without access to an external knowledge source. Our RAG models achieve state-of-the-art results on open Natural Questions [29], WebQuestions [3] and CuratedTrec [2] and strongly outperform recent approaches that use specialised pre-training objectives on TriviaQA [24]. Despite these being extractive tasks, we find that unconstrained generation outperforms previous extractive approaches. For knowledge-intensive generation, we experiment with MS-MARCO [1] and Jeopardy question generation, and we find that our models generate responses that are more factual, specific, and diverse than a BART baseline. For FEVER [56] fact verification, we achieve results within 4.3% of state-of-the-art pipeline models which use strong retrieval supervision. Finally, we demonstrate that the non-parametric memory can be replaced to update the models’ knowledge as the world changes. [^1] Code to run experiments with RAG has been open-sourced as part of the HuggingFace Transformers Library [66] and can be found at https://github.com/huggingface/transformers/blob/master/examples/rag/. An interactive demo of RAG models can be found at https://huggingface.co/rag/
@@
Figure 1: Overview of our approach. We combine a pre-trained retriever (Query Encoder + Document Index) with a pre-trained seq2seq model (Generator) and fine-tune end-to-end. For query $x$, we use Maximum Inner Product Search (MIPS) to find the top-K documents $z_{i}$. For final prediction $y$, we treat $z$ as a latent variable and marginalize over seq2seq predictions given different documents.
@@ 2 Methods
We explore RAG models, which use the input sequence $x$ to retrieve text documents $z$ and use them as additional context when generating the target sequence $y$. As shown in Figure 1, our models leverage two components: (i) a retriever $p_{\eta}(z|x)$ with parameters $\eta$ that returns (top-K truncated) distributions over text passages given a query $x$ and (ii) a generator $p_{\theta}(y_{i}|x,z,y_{1:i-1})$ parametrized by $\theta$ that generates a current token based on a context of the previous i1i-1 tokens $y_{1:i-1}$, the original input $x$ and a retrieved passage $z$.
To train the retriever and generator end-to-end, we treat the retrieved document as a latent variable. We propose two models that marginalize over the latent documents in different ways to produce a distribution over generated text. In one approach, RAG-Sequence, the model uses the same document to predict each target token. The second approach, RAG-Token, can predict each target token based on a different document. In the following, we formally introduce both models and then describe the $p_{\eta}$ and $p_{\theta}$ components, as well as the training and decoding procedure.
@@ 2.1 Models
The RAG-Sequence model uses the same retrieved document to generate the complete sequence. Technically, it treats the retrieved document as a single latent variable that is marginalized to get the seq2seq probability $p(y|x)$ via a top-K approximation. Concretely, the top K documents are retrieved using the retriever, and the generator produces the output sequence probability for each document, which are then marginalized,
$p_{\text{\tiny{RAG-Sequence}}}(y|x)\;\approx\;\sum_{\mathclap{z\in\text{top-}k(p(\cdot|x))}}p_{\eta}(z|x)p_{\theta}(y|x,z)\;=\;\sum_{\mathclap{z\in\text{top-}k(p(\cdot|x))}}p_{\eta}(z|x)\prod_{i}^{N}p_{\theta}(y_{i}|x,z,y_{1:i-1})$
@@
In the RAG-Token model we can draw a different latent document for each target token and marginalize accordingly. This allows the generator to choose content from several documents when producing an answer. Concretely, the top K documents are retrieved using the retriever, and then the generator produces a distribution for the next output token for each document, before marginalizing, and repeating the process with the following output token, Formally, we define:
$p_{\text{\tiny{RAG-Token}}}(y|x)\;\approx\;\prod_{i}^{N}\;\sum_{z\in\text{top-}k(p(\cdot|x))}p_{\eta}(z|x)p_{\theta}(y_{i}|x,z_{i},y_{1:i-1})$
@@ 2.2 Retriever: DPR
$p_{\eta}(z|x)\propto\exp\left(\mathbf{d}(z)^{\top}\mathbf{q}(x)\right)\;\;\;\;\;\;\;\;\;\mathbf{d}(z)=\text{BERT}_{d}(z),\;\;\mathbf{q}(x)=\text{BERT}_{q}(x)$
where $\mathbf{d}(z)$ is a dense representation of a document produced by a BERTBASE document encoder [8], and $\mathbf{q}(x)$ a query representation produced by a query encoder, also based on BERTBASE. Calculating $\text{top-k}(p_{\eta}(\cdot|x))$, the list of $k$ documents $z$ with highest prior probability $p_{\eta}(z|x)$, is a Maximum Inner Product Search (MIPS) problem, which can be approximately solved in sub-linear time [23]. We use a pre-trained bi-encoder from DPR to initialize our retriever and to build the document index. This retriever was trained to retrieve documents which contain answers to TriviaQA [24] questions and Natural Questions [29]. We refer to the document index as the non-parametric memory.
@@ 2.3 Generator: BART
The generator component $p_{\theta}(y_{i}|x,z,y_{1:i-1})$ could be modelled using any encoder-decoder. We use BART-large [32], a pre-trained seq2seq transformer [58] with 400M parameters. To combine the input $x$ with the retrieved content $z$ when generating from BART, we simply concatenate them. BART was pre-trained using a denoising objective and a variety of different noising functions. It has obtained state-of-the-art results on a diverse set of generation tasks and outperforms comparably-sized T5 models [32]. We refer to the BART generator parameters $\theta$ as the parametric memory henceforth.
@@ 2.4 Training
We jointly train the retriever and generator components without any direct supervision on what document should be retrieved. Given a fine-tuning training corpus of input/output pairs $(x_{j},y_{j})$, we minimize the negative marginal log-likelihood of each target, $\sum_{j}-\log p(y_{j}|x_{j})$ using stochastic gradient descent with Adam [28]. Updating the document encoder $\text{BERT}_{d}$ during training is costly as it requires the document index to be periodically updated as REALM does during pre-training [20]. We do not find this step necessary for strong performance, and keep the document encoder (and index) fixed, only fine-tuning the query encoder $\text{BERT}_{q}$ and the BART generator.
@@ 2.5 Decoding
At test time, RAG-Sequence and RAG-Token require different ways to approximate $\operatorname*{arg\,max}_{y}p(y|x)$.
@@
The RAG-Token model can be seen as a standard, autoregressive seq2seq generator with transition probability: $p^{\prime}_{\theta}(y_{i}|x,y_{1:i-1})=\sum_{z\in\text{top-}k(p(\cdot|x))}p_{\eta}(z_{i}|x)p_{\theta}(y_{i}|x,z_{i},y_{1:i-1})$ To decode, we can plug $p^{\prime}_{\theta}(y_{i}|x,y_{1:i-1})$ into a standard beam decoder.
@@
For RAG-Sequence, the likelihood $p(y|x)$ does not break into a conventional per- token likelihood, hence we cannot solve it with a single beam search. Instead, we run beam search for each document $z$, scoring each hypothesis using $p_{\theta}(y_{i}|x,z,y_{1:i-1})$. This yields a set of hypotheses $Y$, some of which may not have appeared in the beams of all documents. To estimate the probability of an hypothesis $y$ we run an additional forward pass for each document $z$ for which $y$ does not appear in the beam, multiply generator probability with $p_{\eta}(z|x)$ and then sum the probabilities across beams for the marginals. We refer to this decoding procedure as “Thorough Decoding.” For longer output sequences, $|Y|$ can become large, requiring many forward passes. For more efficient decoding, we can make a further approximation that $p_{\theta}(y|x,z_{i})\approx 0$ where $y$ was not generated during beam search from $x,z_{i}$. This avoids the need to run additional forward passes once the candidate set $Y$ has been generated. We refer to this decoding procedure as “Fast Decoding.”
@@ 3 Experiments
We experiment with RAG in a wide range of knowledge-intensive tasks. For all experiments, we use a single Wikipedia dump for our non-parametric knowledge source. Following Lee et al. [31] and Karpukhin et al. [26], we use the December 2018 dump. Each Wikipedia article is split into disjoint 100-word chunks, to make a total of 21M documents. We use the document encoder to compute an embedding for each document, and build a single MIPS index using FAISS [23] with a Hierarchical Navigable Small World approximation for fast retrieval [37]. During training, we retrieve the top $k$ documents for each query. We consider $k\in\{5,10\}$ for training and set $k$ for test time using dev data. We now discuss experimental details for each task.
@@ 3.1 Open-domain Question Answering
Open-domain question answering (QA) is an important real-world application and common testbed for knowledge-intensive tasks [20]. We treat questions and answers as input-output text pairs $(x,y)$ and train RAG by directly minimizing the negative log-likelihood of answers. We compare RAG to the popular extractive QA paradigm [5, 7, 31, 26], where answers are extracted spans from retrieved documents, relying primarily on non-parametric knowledge. We also compare to “Closed-Book QA” approaches [52], which, like RAG, generate answers, but which do not exploit retrieval, instead relying purely on parametric knowledge. We consider four popular open-domain QA datasets: Natural Questions (NQ) [29], TriviaQA (TQA) [24]. WebQuestions (WQ) [3] and CuratedTrec (CT) [2]. As CT and WQ are small, we follow DPR [26] by initializing CT and WQ models with our NQ RAG model. We use the same train/dev/test splits as prior work [31, 26] and report Exact Match (EM) scores. For TQA, to compare with T5 [52], we also evaluate on the TQA Wiki test set.